OpenStudy (anonymous):

What is the sum of the geometric sequence –4, 24, –144, … if there are 8 terms?

OpenStudy (anonymous):

the common ratio r = -24/4 = -6 (and 144 / 24 = -6) sum = (-4) * (1 -(-6)^8) -------- 1 - (-6)

OpenStudy (anonymous):

can u finish this off?

OpenStudy (anonymous):

i dont think im getting the right naswer

OpenStudy (anonymous):

you should get 959780

OpenStudy (anonymous):

positive?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

aw thanks! :)

OpenStudy (anonymous):

welcome